# Channel Estimation for Beyond Diagonal RIS via Tensor Decomposition

### v1

@@ VI-A BD-RIS design for the BTKF algorithm
The first step of the BTKF algorithm (see Figure 3) corresponds to filtering the received pilot tensor by the compact 3-mode unfolding $\bm{S}_{3}=\bm{\Theta}\otimes[\bm{\mathcal{Z}}]_{(3)}$ of the BD-RIS tensor. This step is the same as that of the traditional LS estimation method and is subject to the same requirements.
If $K\geq\bar{N}^{2}Q$, or, equivalently, if $K_{1}=\bar{N}^{2}$ and $K_{2}\geq Q$, then $\bm{S}_{3}$ is column-orthogonal, which implies $\bm{S}^{\text{H}}_{3}\bm{S}_{3}=(K/\bar{N})\bm{I}_{\bar{N}^{2}Q}$. This property can be checked as follows. We can first note that $K\geq\bar{N}^{2}Q$ implies $K_{1}=\bar{N}^{2}$ and $K_{2}=K/\bar{N}^{2}$. In this case, the 3-mode unfolding of the circulant tensor $\bm{\mathcal{Z}}\in\mathbb{C}^{\bar{N}\times\bar{N}\times\bar{N}^{2}}$ is a square $\bar{N}^{2}\times\bar{N}^{2}$ circulant matrix with orthogonal rows constructed from circular shifts of the DFT basis in vectorized form, which implies $[\bm{\mathcal{Z}}]^{\text{H}}_{(3)}[\bm{\mathcal{Z}}]_{(3)}=\bar{N}\bm{I}_{\bar{N}^{2}}$. From (68), we have $[\bm{\mathcal{S}}^{(q)}]^{\text{H}}_{(3)}[\bm{\mathcal{S}}^{(q)}]_{(3)}=(\bm{\Theta}^{\text{H}}\bm{\Theta})\otimes([\bm{\mathcal{Z}}]^{\text{H}}_{(3)}[\bm{\mathcal{Z}}]_{(3)})=K/\bar{N}^{2}\otimes\bar{N}\bm{I}_{\bar{N}^{2}}=(K/\bar{N})\bm{I}_{\bar{N}^{2}}$. Hence, it follows that $\bm{S}^{\text{H}}_{3}\bm{S}_{3}=\bm{I}_{\bar{N}^{2}Q}$.
Due to this column-orthogonality property, a simple matched filtering can replace the pseudo-inverse in step 3 of Algorithm 1 while offering optimized performance [5].
@@ VI-B BD-RIS design for the BTALS algorithm
The BTALS algorithm can provide unique estimates of the individual channel matrices by capitalizing on the essential uniqueness properties of block Tucker decompositions [24], [33]. As discussed in section V-D, in our context, the uniqueness of our BD-RIS block Tucker model (28) assumes that the BD-RIS training tensor does not have proportional frontal slices, which means no slices in the set $\{\bm{\mathcal{S}}^{(q)}_{..1},\ldots,\bm{\mathcal{S}}^{(q)}_{..K}\}$, $\forall q\in\{1,\ldots,Q\}$ are scalar multiples of each other. To cope with this assumption while preserving structural properties, we design the BD-RIS training tensor as
@@
where $\bar{\bm{\mathcal{S}}}^{(q)}_{..k}$ follows the design in (65), and $\bm{W}^{(q)}\doteq[\bm{w}^{(q)}_{1},\ldots,\bm{w}^{(q)}_{K}]^{\text{T}}\in\mathbb{C}^{K\times\bar{N}}$ is a matrix of exponentials, whose $k$-th row is given by $\bm{w}^{(q)}_{k}\doteq[1\,\,e^{j\psi^{(q)}_{k,1}},\ldots,e^{j\psi^{(q)}_{k,\bar{N}-1}}]^{\text{T}}$, with phases $\{\psi^{(q)}_{k,n}\}$ uniformly distributed between $[0,2\pi)$, $k=1,\ldots,K$, $q=1,\ldots,Q$.
The random phase shifts generated by the transformations in (69) ensure that the $K$ frontal slices of the BD-RIS training tensor are uncorrelated. This eliminates proportional slices that may result from the design in (65) depending on the values assumed for $\bar{N}$ and $Q$, ensuring the uniqueness of the channel estimates. Note also that the random left/right rotations applied to the basis tensor preserve the orthogonality of the frontal slices, i.e., $\bm{\mathcal{S}}^{(q)}_{..k}\bm{\mathcal{S}}^{(q)\text{H}}_{..k}=\bm{\mathcal{S}}^{(q)\text{H}}_{..k}\bm{\mathcal{S}}^{(q)}_{..k}=\bm{I}_{\bar{N}}$, $k=1,\ldots,K$.
@@ VII Numerical results
We evaluate the performance of the proposed BTKF and BTALS receivers under different system setups. We also discuss the tradeoffs involving the proposed channel estimation methods and their superior performance compared to the LS scheme [5]. We consider the NMSE of the combined channel $\bm{C}=\bm{H}\hskip 1.42271pt|\otimes|\hskip 1.42271pt\bm{G}\in\mathbb{C}^{M_{R}M_{T}\times\bar{N}^{2}Q}$. The NMSE is given by
$\text{NMSE}\big{(}\hat{\bm{C}}\big{)}=\frac{\left\|\bm{C}-\hat{\bm{C}}\right\|_{\rm F}^{2}}{\left\|\bm{C}\right\|_{\rm F}^{2}}.$
As in [5], we assume i.i.d. channels, and we compare the performance of the different methods for a variety of parameter setups that are shown in each figure.
In Figures 5 to 7, we evaluate the performance of the proposed receiver BTKF with the baseline LS of [5]. It is worth mentioning that the BTALS is omitted here due to the fact that it achieves the same performance as the BTKF in the considered scenarios. Since the BTALS is the only receiver that can operate at a lower training overhead, we will discuss its performance in a separate section (Figure 8 to 10). For all results, orthogonal pilot sequences are assumed, and their length is fixed to its minimum value $T_{\text{min}}=M_{T}$.
@@
In Figure 5, we compare the NMSE of the combined channel by fixing $K=256$, which is the minimum value ($K_{\text{min}}=\bar{N}^{2}Q$) for the configuration with largest group size $\bar{N}=4$. First, we observe that BTKF offers higher estimation accuracies than the baseline LS method, regardless of the group size. Indeed, the LS method estimates the combined channel as a classical MIMO channel estimation problem, which is blind to its inherent Kronecker factorization structure. Thus, the performance of the reference method is sensitive to the number of groups and degrades as the group size ($\bar{N}$) increases, while the BTKF method is insensitive to it. Additionally, we can see that the performance gains of BTKF over LS increase with $\bar{N}$. Such a gain comes from the channel separation property of BTKF that exploits the Kronecker structure of the combined channel.
@@
In Figure 6, we assume the minimum training overhead for each considered BD-RIS configuration, i.e., $K_{\text{min}}=\bar{N}^{2}Q$ for each setup, that ensure a unique solution. It is interesting to note that the baseline LS scheme [5] has the same performance in all configurations. Hence, the total number of channel coefficients to be estimated is the same in all the training configurations considered in this figure. In contrast, the BTKF method estimates fewer channel coefficients than LS. The gap in terms of the number of channel coefficients to be estimated increases as the group size, which explains the performance gap between the two methods as $\bar{N}$ increases. Note that, in the single-connected case ($\bar{N}=1$), the numbers of estimated channel coefficients are the same for LS and BTKF. However, even in this case, the BTKF still offers a performance gain over LS. From this set of results, one can conclude that as the group size increases, the noise rejection gain offered by the rank-one approximation step of BTKF also increases. In contrast, the LS method has nearly the same performance regardless of the group size since it does not take advantage of the Kronecker structure of the combined channel.
In Figure 7, we fix the group size $\bar{N}=4$ and the total number of groups $Q=16$, implying $N=64$ RIS elements. In this experiment, we compare the performance of BTKF and LS for different numbers of antennas $M_{T}$ and $M_{R}$ at the transmitter and receiver, respectively. Note that the pilot sequence length is adjusted in each configuration to its minimum value ($T=M_{T}$). We observe that the performance of the LS estimator is the same regardless of the number of transmit/receive antennas. Indeed, although the pilot length increases to ensure orthogonality, the number of channel coefficients of the combined channel also increases with $M_{T}$ and $M_{R}$. This means that the LS estimator cannot benefit from the added spatial degrees of freedom to improve channel estimation performance. In contrast, the BTKF method efficiently benefits from more transmit and receive spatial degrees of freedom by capitalizing on the Kronecker structure of the combined channel. Such a gain comes from the noise rejection gains provided by the rank-one approximation problems in (56), whose dimensions increase with $M_{T}$ and $M_{R}$, yielding more accurate estimates of $\bm{G}$ and $\bm{H}$ and, consequently, of the combined channel $\bm{C}=\hat{\bm{H}}\hskip 1.42271pt|\otimes|\hskip 1.42271pt\hat{\bm{G}}$.
@@
In the next experiments (Figures 8 to 10), we turn our attention to the BTALS algorithm and study its performance for several system configurations. The focus is on very low training overhead setups, where $K<<\bar{N}^{2}Q$. Hence, these figures do not show the results of the LS and BTKF methods since they cannot operate in the considered challenging setups over the full range of values of $K$ due to their more restrictive requirements on this training parameter.
@@
In Figure 9, we show the performance of BTALS as a function of the number $Q$ of groups, going from the fully-connected case ($Q=1$) to the single-connected case ($Q=64$). We assume $N=\bar{N}Q=64$ RIS elements, $K=64$ blocks, and an SNR of $20$ dB while considering configurations with different numbers of transmit and receive antennas. As expected, the performance increases as more transmit/receive antennas are used, showing the effectiveness of BTALS in converting spatial diversity gains at both link ends into more accurate channel estimates despite the increase in the number of estimated channel coefficients. These results corroborate the gains of tensor-based processing for BD-RIS channel estimation. We can also see that when the number of antennas and the transmit/receive increases, the gap among the single-, group- and fully-connected architectures reduces.
Figure 11: Average number of iterations for BTALS convergence as a function of the number $K$ of training blocks.
@@
In Figure 11, we study the convergence performance of BTALS. First, we can see that lower training lengths imply more iterations to achieve convergence. Indeed, increased time diversity is available when more training blocks are used, which helps BTALS to speed up its iterative process. For very low training lengths ($K\leq 32$), the required number of iterations is more sensitive to the group size $\bar{N}$. However, for most of the considered values of $K$, the convergence behavior is similar regardless of the chosen group-connected configuration. It can also be noted that the single-connected case using optimal DFT design for the RIS training matrix needs fewer iterations than the group-connected cases, although the gap is not significant.
@@
In Figure 12, we compare LS, BTKF, and BTALS in terms of computational complexity. We plot the complexity order of each algorithm as a function of the group size $\bar{N}$ for a BD-RIS with $N=64$ elements. Recall that $\bar{N}=1$ corresponds to the single-connected case, while $\bar{N}=64$ to the fully-connected case. As expected, the computational complexity increases with the group size in all cases. Moreover, since the convergence of BTALS is sensitive to the SNR (see Fig. 11), its overall complexity will increase at low SNRs. This is not the case with BTKF, whose complexity does not depend on the SNR due to its closed-form nature. Such a complexity gap is the price BTALS pays to operate at a very low overhead compared to the BTFK and the LS methods. In addition, BTKF and LS have comparable complexities in this scenario since the complexity is dominated by the matched filtering step (the same for both LS and BTKF).

### v2

@@ VI-A BD-RIS design for the BTKF algorithm
The initial step of the BTKF algorithm involves left-filtering the 3-mode unfolding $[\bm{\mathcal{Y}}]_{(3)}$ using the compact 3-mode unfolding $\bm{S}_{3}=\bm{\Theta}\otimes[\bm{\mathcal{Z}}]_{(3)}$ of the BD-RIS training tensor.
Property 1: If $K_{1}=\bar{N}^{2}$ and $K_{2}\geq Q$, then $\bm{\Theta}$ and $[\bm{\mathcal{Z}}]_{(3)}$ are column-orthogonal matrices such that $\bm{S}_{3}$ has full column rank and column-orthogonal, satisfying $\bm{S}^{\text{H}}_{3}\bm{S}_{3}=(K/\bar{N})\bm{I}_{\bar{N}^{2}Q}$.
This property can be checked as follows. We first note that $K\geq\bar{N}^{2}Q$ implies $K_{1}=\bar{N}^{2}$ and $K_{2}=K/\bar{N}^{2}$. In this case, $[\bm{\mathcal{Z}}]_{(3)}$ is a square $\bar{N}^{2}\times\bar{N}^{2}$ circulant matrix with orthogonal rows constructed from circular shifts of the DFT basis in vectorized form, which implies $[\bm{\mathcal{Z}}]^{\text{H}}_{(3)}[\bm{\mathcal{Z}}]_{(3)}=\bar{N}\bm{I}_{\bar{N}^{2}}$. From (68), we have $[\bm{\mathcal{S}}^{(q)}]^{\text{H}}_{(3)}[\bm{\mathcal{S}}^{(q)}]_{(3)}=(\bm{\Theta}^{\text{H}}\bm{\Theta})\otimes([\bm{\mathcal{Z}}]^{\text{H}}_{(3)}[\bm{\mathcal{Z}}]_{(3)})=K/\bar{N}^{2}\otimes\bar{N}\bm{I}_{\bar{N}^{2}}=(K/\bar{N})\bm{I}_{\bar{N}^{2}}$. Hence, it follows that $\bm{S}^{\text{H}}_{3}\bm{S}_{3}=\bm{I}_{\bar{N}^{2}Q}$. Due to this column-orthogonality property, a simple matched filtering can replace the pseudo-inverse in step 3 of Algorithm 1 while offering optimum performance [5].
@@ VI-B BD-RIS design for the BTALS algorithm
The BTALS algorithm estimates the individual channel matrices from the 1-mode and 2-mode unfoldings $[\bm{\mathcal{Y}}]_{(1)}$ and $[\bm{\mathcal{Y}}]_{(2)}$ according to (62).
Property 2: If no proportional columns exist in the set $\{\bm{\mathcal{S}}^{(q)}_{..1},\ldots,\bm{\mathcal{S}}^{(q)}_{..K}\}$, $\forall q\in\{1,\ldots,Q\}$, meaning that the BD-RIS training tensor does not have proportional frontal slices, then $\bm{S}_{1}$ and $\bm{S}_{2}$ have full row-rank and are row-orthogonal, satisfying $\bm{S}^{\text{H}}_{1}\bm{S}_{1}=\bm{S}^{\text{H}}_{2}\bm{S}_{2}=K\bm{I}_{\bar{N}Q}$.
To verify this property, first note that according to (64) and (65), the $k$-th frontal slice $\bm{\mathcal{S}}^{(q)}_{..k}$ is orthogonal by construction, i.e., $\bm{\mathcal{S}}^{(q)}_{..k}\bm{\mathcal{S}}^{(q)\text{H}}_{..k}=\bm{\mathcal{S}}^{(q)\text{H}}_{..k}\bm{\mathcal{S}}^{(q)}_{..k}=\bm{I}_{\bar{N}}$, $k=1,\ldots,K$. Since $\bm{S}_{1}$ (resp. $\bm{S}_{2}$) concatenates the columns (resp. rows) of the $K$ slices $\{\bm{\mathcal{S}}^{(q)}_{..k}\}$ (see (29)-(30)), they will have full rank $\bar{N}Q$ if their columns (resp. rows) are non-proportional, which implies $\bm{\mathcal{S}}_{..k}\neq\alpha\bm{\mathcal{S}}_{..k^{\prime}}$, $\forall k\neq k^{\prime}$, $k,k^{\prime}\in\{1,\ldots,K\}$, where $\alpha$ is a scalar. The constraint $\bm{S}^{\text{H}}_{1}\bm{S}_{1}=\bm{S}^{\text{H}}_{2}\bm{S}_{2}=K\bm{I}_{\bar{N}Q}$ follows from the orthogonality of each matrix $\bm{\mathcal{S}}^{(q)}_{..k}$, $k=1,\ldots,K$.
To ensure Property 2 while preserving its structural properties, the BD-RIS training tensor is designed as
@@
where $\bar{\bm{\mathcal{S}}}^{(q)}_{..k}$ is given by (65) and $\bm{W}^{(q)}\doteq[\bm{w}^{(q)}_{1},\ldots,\bm{w}^{(q)}_{K}]^{\text{T}}\in\mathbb{C}^{K\times\bar{N}}$ is a matrix of exponentials, whose $k$-th row is given by $\bm{w}^{(q)}_{k}\doteq[1\,\,e^{j\psi^{(q)}_{k,1}},\ldots,e^{j\psi^{(q)}_{k,\bar{N}-1}}]^{\text{T}}$, where $\{\psi^{(q)}_{k,n}\}$ are uniformly distributed between $[0,2\pi)$, $k=1,\ldots,K$, $q=1,\ldots,Q$. The random phase shifts in (69) ensure that the $K$ frontal slices of the BD-RIS training tensor are non-proportional, regardless of the values of $\bar{N}$ and $Q$. Note also that the orthogonality of the frontal slices are preserved after these random phase shifts, i.e., $\bm{\mathcal{S}}^{(q)}_{..k}\bm{\mathcal{S}}^{(q)\text{H}}_{..k}=\bm{\mathcal{S}}^{(q)\text{H}}_{..k}\bm{\mathcal{S}}^{(q)}_{..k}=\bm{I}_{\bar{N}}$, $k=1,\ldots,K$.
VI-C Discussion
It is clear that the BTALS algorithm has less restrictive requirements on the training length $K$ than the BTKF algorithm. This comes from the fact that the latter only exploits the 3-mode unfolding of the received pilot tensor (as is the case with the LS method [5]), and channel decoupling is carried out after the LS filtering step, which is the bottleneck restricting the training length. On the other hand, the BTALS algorithm exploits the 1-mode and 2-mode unfoldings of the received pilot tensor to estimate the involved channel matrices directly. The associated LS steps involve more relaxed requirements for the training length $K$ compared to BTKF.
It is worth mentioning that the BTFK method accomplishes channel estimation in closed form utilizing $Q$ independent and parallel processing routines, which implies a shorter processing delay than BTALS. The latter consists of a sequential process of alternating estimation steps, where the channel matrices associated with the $Q$ BD-RIS groups are jointly estimated. Therefore, there is a longer processing delay than BTKF. From a practical perspective, optimized procedures for computing rank-one approximations and the possibility of parallel processing are attractive features of BTKF despite its higher training overhead than BTALS. Thus, there is a clear trade-off between both algorithms involving complexity, training overhead, and processing delay.
@@ VII Numerical results
We evaluate the performance of the proposed BTKF and BTALS receivers under different system setups. We also discuss the tradeoffs involving the proposed channel estimation methods and their superior performance compared to the LS scheme [5]. We consider the NMSE of the composite channel $\bm{T}=\bm{H}\hskip 1.42271pt|\otimes|\hskip 1.42271pt\bm{G}\in\mathbb{C}^{M_{R}M_{T}\times\bar{N}^{2}Q}$. The NMSE is given by
$\text{NMSE}\big{(}\hat{\bm{T}}\big{)}=\frac{\left\|\bm{T}-\hat{\bm{T}}\right\|_{\rm F}^{2}}{\left\|\bm{T}\right\|_{\rm F}^{2}}.$
As in [5], we assume i.i.d. channels, and we compare the performance of the different methods for various parameter setups shown in each figure.
In Figures 3 to 5, we evaluate the performance of the proposed receiver BTKF with the baseline LS of [5]. It is worth mentioning that the BTALS is omitted here due to the fact that it achieves the same performance as the BTKF in the considered scenarios. Since the BTALS is the only receiver that can operate at a lower training overhead, we will discuss its performance in a separate section (Figure 6 to 13). For all results, orthogonal pilot sequences are assumed, and their length is fixed to its minimum value $T_{\text{min}}=M_{T}$.
@@
In Figure 3, we compare the NMSE of the composite channel by fixing $K=256$, which is the minimum value ($K_{\text{min}}=\bar{N}^{2}Q$) for the configuration with largest group size $\bar{N}=4$. First, BTKF offers higher estimation accuracies than the baseline LS method, regardless of the group size. Indeed, conventional LS estimates the composite channel as a classical MIMO channel estimation problem, which is blind to its inherent Kronecker structure. Thus, the performance of the reference method is sensitive to the number of groups and degrades as the group size ($\bar{N}$) increases, while the BTKF method is insensitive to it. Additionally, we can see that the performance gains of BTKF over LS increase with $\bar{N}$. Such a gain comes from the channel separation property of BTKF that exploits the Kronecker structure of the composite channel.
@@
In Figure 4, we assume the minimum training overhead for each considered BD-RIS configuration, i.e., $K_{\text{min}}=\bar{N}^{2}Q$ for each setup that ensures a unique solution. It is interesting to note that the baseline LS scheme [5] has the same performance in all configurations. Hence, the estimated number of channel coefficients is the same in all the configurations considered in this figure. In contrast, the BTKF method estimates fewer channel coefficients than LS. The gap in the number of channel coefficients to be estimated increases as the group size increases, which explains the performance gap between the two methods as $\bar{N}$ increases. Note that, in the single-connected case ($\bar{N}=1$), the estimated channel coefficients are the same for LS and BTKF. However, the BTKF still offers a performance gain over LS, even in this case. From this set of results, one can conclude that as the group size increases, the noise rejection gain offered by the rank-one approximation step of BTKF also increases. In contrast, the LS method has nearly the same performance regardless of the group size since it does not take advantage of the Kronecker structure of the composite channel.
In Figure 5, we fix the group size $\bar{N}=4$ and the total number of groups $Q=16$, implying $N=64$ RIS elements. In this experiment, we compare the performance of BTKF and LS for different numbers of antennas $M_{T}$ and $M_{R}$ at the transmitter and receiver, respectively. Note that the pilot sequence length is adjusted in each configuration to its minimum value ($T=M_{T}$). We observe that the performance of the LS estimator is the same regardless of the number of transmit/receive antennas. Indeed, although the pilot length increases to ensure orthogonality, the number of composite channel coefficients also increases with $M_{T}$ and $M_{R}$. This means the LS estimator cannot benefit from the added spatial degrees of freedom to improve channel estimation performance. In contrast, the BTKF method efficiently benefits from more transmit and receive spatial degrees of freedom by capitalizing on the Kronecker structure of the composite channel. Such a gain comes from the noise rejection gains provided by the rank-one approximation problems in (51), whose dimensions increase with $M_{T}$ and $M_{R}$, yielding more accurate estimates of $\bm{G}$ and $\bm{H}$ and, consequently, of the composite channel $\hat{\bm{T}}=\hat{\bm{H}}\hskip 1.42271pt|\otimes|\hskip 1.42271pt\hat{\bm{G}}$.
@@
In the following experiments (Figures 6 to 13), we turn our attention to the BTALS algorithm and study its performance for several system configurations. The focus is on low training overhead setups, where $K<<\bar{N}^{2}Q$. Hence, these figures do not show the results of the LS and BTKF methods since they cannot operate in the considered challenging setups over the full range of values of $K$ due to their more restrictive requirements on this training parameter.
@@
In Figure 7, we show the performance of BTALS as a function of the number $Q$ of groups, going from the fully-connected case ($Q=1$) to the single-connected case ($Q=64$). We assume $N=\bar{N}Q=64$ RIS elements, $K=64$ blocks, and an SNR of $20$ dB while considering configurations with different numbers of transmit and receive antennas. As expected, the performance increases as more transmit/receive antennas are used, showing the effectiveness of BTALS in converting spatial diversity gains at both link ends into more accurate channel estimates despite the increase in the number of estimated channel coefficients. These results corroborate the gains of tensor-based processing for BD-RIS channel estimation. We can also see that when the number of antennas and the transmit/receive increases, the gap among the single-, group-, and fully-connected architectures reduces.
Figure 9: Reconstruction error $\epsilon$ per iteration for different setups.
Figure 10: Number of iterations to convergence vs. SNR.
@@
In Figures 9 and 10, we evaluate the convergence properties of the BTALS algorithm for different system setups. Figure 9 corroborates the monotonic convergence of BTALS for all BD-RIS configurations, i.e., from the single-connected case ($\bar{N}=1$) to the fully-connected case ($\bar{N}=64$), in the noiseless case. We can also observe that, regardless of the specific setup, the BTALS algorithm always converges to the machine’s precision within 40-50 iterations. Figure 10 shows a different view of the convergence by depicting the required number of iterations to BTALS convergence for different system setups. As expected, the convergence is faster as the SNR increases. In particular, for a medium SNR above $10$ dB, convergence is always achieved with less than 30 iterations. It is worth mentioning that these results assume only $K=40$ blocks, which is significantly smaller than $\bar{N}^{2}Q$.
@@
In Figure 11, we compare LS, BTKF, and BTALS in terms of computational complexity. We plot the complexity order of each algorithm as a function of the group size $\bar{N}$ for a BD-RIS with $N=64$ elements. Recall that $\bar{N}=1$ corresponds to the single-connected case, while $\bar{N}=64$ corresponds to the fully-connected case. As expected, the computational complexity increases with the group size in all cases. Moreover, since the convergence of BTALS is sensitive to the SNR (see Fig. 10), its overall complexity will increase at low SNRs. This is not the case with BTKF, whose complexity does not depend on the SNR due to its closed-form nature. Such a complexity gap is the price BTALS pays to operate at a very low overhead compared to the BTFK and the LS methods. In addition, BTKF and LS have comparable complexities in this scenario since the complexity is dominated by the matched filtering step (the same for both LS and BTKF).
Figure 12: BTALS vs. SNR performance for low-rank channel $\bm{H}$, considering different number of channel paths.
Figure 13: BTALS performance for low-rank channel $\bm{H}$ as a function of the number of channel paths and SNR of $25$ dB.
Our proposed algorithms can also operate under spatially correlated channels or scattering-poor propagation (low-rank) channels. In Figure 12, we show the NMSE performance of the proposed BTALS algorithm under multipath (geometric) channels. More specifically, we assume that the channel between the transmitter and the BD-RIS ($\bm{H}$) is modeled as
$\bm{H}=\sum\limits_{l=1}^{L}\bm{a}^{(\text{Tx})}_{l}\gamma_{l}\bm{a}^{(\text{RIS})\text{T}}_{l}\in\mathbb{C}^{M_{T}\times N},$
where $\bm{a}^{(\text{RIS})}_{l}\in\mathbb{C}^{N\times 1}$ and $\bm{a}^{(\text{Tx})}_{l}\in\mathbb{C}^{M_{T}\times 1}$ are the steering vectors associated with the arrival and departure angles at the BD-RIS and Tx, respectively. The angles are randomly generated between $[-\pi,\pi]$ according to a uniform distribution, while the $l$-th path amplitude $\gamma_{l}$ is a zero mean unit variance complex Gaussian random variable, $l=1,\ldots,L$. The channel matrix $\bm{G}$ linking the BD-RIS to the receiver follows a Rayleigh fading model, corresponding to a rich-scattering propagation. Considering the rank-deficient channel case, improved performance is obtained when the number of channel paths linking the transmitter to the BD-RIS increases. This is expected since, in this case, the rank of $\bm{H}$ corresponds to the number $L$ of paths, where $L<\text{min}\{M_{T},N\}$. The higher the $L$, the better the condition number of the channel matrix $\bm{H}$. Note, however, that the performance gap is not significant.
Figure 13 shows the performance of the BTALS as a function of the number of paths of $\bm{H}$ for different values of $K$ and an SNR of 25dB. The results show that poor scattering scenarios with few paths become more sensitive to the training length $K$. As the number of channel paths increases, the choice of this parameter has a negligible influence on performance.
The BTKF and BTALS algorithms have interesting tradeoffs. When training overhead is not critical, BTKF is preferable due to its lower complexity, closed-form solution, and small processing delay, since each rank-one factorization is independent and can be computed in parallel. On the other hand, when minimizing the training overhead is crucial (especially for strongly connected BD-RIS), BTALS is more attractive due to its flexibility in operating with considerably fewer training blocks than the LS and BTKF methods with no performance degradation.